# Plurals: A System for Guiding LLMs Via Simulated Social Ensembles

### v5

@@ Abstract.
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: ; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Computing methodologies Artificial intelligenceccs: Computing methodologies Multi-agent systemsccs: Computing methodologies Intelligent agentsccs: Human-centered computing Interaction paradigmsccs: Human-centered computing Interactive systems and toolsccs: Human-centered computing Open source softwareccs: Human-centered computing Interaction design theory, concepts and paradigms
@@ 1. Introduction
Figure 1. System diagram of Plurals—an end-to-end generator of simulated social ensembles. (1) Agents complete tasks within (2) Structures, with communication optionally summarized by (3) Moderators. Plurals integrates with government datasets (1a) and templates inspired by democratic deliberation theory (1b). The building block is Agents, which are large language models (LLMs) that have system instructions and tasks. System instructions can be generated from user input, government datasets (American National Election Studies; ANES), and templates from deliberative democracy literature (bachtiger_deliberative_2018, ). Agents exist within Structures, which define what information is shared. Combination instructions tell Agents how to combine the responses of other Agents when deliberating in the Structure. Users can customize an Agent’s combination instructions or use existing templates drawn from deliberation literature and beyond. Moderators aggregate responses from multi-agent deliberation.
@@
System Diagram
There is a fundamental tension between how generative AI models are built and how they are used. Companies typically build a small number of foundation or “generalist” models that dominate the market (vipra_market_2023, ). However, these generalist models are used by a diverse base of users—with varying preferences and values. Invariably, this tension sparked allegations of bias, with supposedly neutral models accused of favoring certain viewpoints (feng_pretraining_2023, ; braun_politics_nodate, ; fujimoto_revisiting_2023, ).
While a tempting solution is to aim for models that have “no bias” and hold a “view from nowhere” (haraway_situated_1997, ), truly neutral models are likely infeasible. Some scholars argue that all knowledge is situated (haraway_situated_1997, ). But with open-ended text generation, defining some unbiased ground truth is especially difficult. For many use cases, there is no unbiased ground truth. This difficulty is compounded by the fact that users can ask models a large variety of questions. Any bias benchmark can only capture an infinitesimal slice of the query space (raji_ai_2021, ).
@@
As an alternative to “bias-free” models, we introduce a new pluralistic AI system (sorensen_roadmap_2024, ), Plurals, that can accomplish these tasks. It is a public-facing Python library (Figure 1 for system overview, Figure 2 for code snippets, see here [^1] https://github.com/josh-ashkinaze/plurals for library). Plurals consists of Agents (optionally integrated with government datasets for nationally representative personas) which deliberate within customizable Structures, with Moderators overseeing deliberation. Plurals is an end-to-end generator of customizable “simulated social ensembles”. We incorporate interaction templates inspired by democratic deliberation theory and integration with government datasets for nationally representative personas. For example, to create an Agent representing a male California resident, our system samples a statistically representative citizen from American National Election Studies, and then uses the citizen’s demographics and political stances as an LLM prompt. We draw on deliberative democracy theory, which emphasizes dialogue between different views (bachtiger_deliberative_2018, ; marti_pluralism_2017, ), as a blueprint. Our work builds on research in deliberation (bachtiger_deliberative_2018, ; morrell_listening_2018, ; brown_deliberation_2018, ; smith_mini-publics_2018, ; fraser_rethinking_1990, ; habermas_structural_1991, ; marti_pluralism_2017, ), pluralistic sociotechnical systems (argyle_out_2023, ; lee_webuildai_2019, ; gordon_jury_2022, ; zhang_policykit_2020, ), and multi-agent AI alignment approaches (vahidov_pluralistic_2004, ; irving_ai_2018, ; talebirad_multi-agent_2023, ; kim_can_2024, ; chan_chateval_2023, ). To our knowledge, Plurals is the first general-purpose toolkit for pluralistic, multi-agent interactions modeled after democratic deliberation.
@@
  • Theoretical: We created a multi-agent system incorporating ideals of democratic deliberation theory. Our system also introduces “interactional pluralism”, a pluralism that exists not only in the distribution of agent properties but also in the protocols governing their interactions. Users can customize how Agents should combine information with each other and the information-sharing structures in which Agents exist.
@@ 1.1. Brief System Overview
Plurals allows users to create simulated social ensembles with Agents, Structures, and Moderators: Agents complete tasks within Structures, which define how information is shared between Agents. Moderators can summarize multi-agent communication. Each abstraction is highly customizable. Agents can use various LLMs and have system instructions set manually, through persona generation, or via American National Election Studies (ANES) integration. Structures vary in information-sharing, complexity, and randomness. For example, users can define custom networks of Agents in a few lines of code (Figure 2). The behavior of Agents within Structures (how they should combine information from other agents) can be tuned via combination instructions. Our package comes pre-populated with templates for personas, combination instructions, and moderators—drawing on democratic deliberation theory and prior work.
@@ 2. System Grounding
Plurals is grounded in deliberation literature, sociotechnical systems that broaden technological perspectives, and multi-agent systems for AI alignment. The result is an end-to-end generator of simulated social ensembles—groups that engage in deliberation. We integrate deliberative theory into our system by incorporating templates of first- and second-generation deliberative ideals and using deliberative theory to inform the structure of AI deliberation. We extend previous work on broadening technological perspectives, such as Argyle et al.’s dataset-based personas (argyle_out_2023, ), Gordon et al.’s “juries” (gordon_jury_2022, ), and Zhang et al.’s PolicyKit (zhang_policykit_2020, ). Our system encompasses individual, group, and governance-level simulations, unlike previous approaches that focused on flexibility at only one of these three levels. By drawing on the concept of deliberative “mini-publics” (groups who engage in deliberation (smith_mini-publics_2018, )), we evolve from aggregative methods (like juries) to a more deliberative approach. Additionally, we contribute to multi-agent AI research by offering a flexible system for creating diverse interaction structures and providing reusable code and shared infrastructure for experiments.
@@ 2.1. Deliberation
Deliberation is defined as “mutual communication that involves weighing and reflecting on preferences, values, and interests regarding matters of common concern” (bachtiger_deliberative_2018, ). As Bächtiger et al. distinguish (bachtiger_deliberative_2018, ), deliberative democracy differs from aggregative democracy. The former centers talking and the latter centers voting—though they can co-occur (e.g., talking before voting (fishkin_consulting_2003, ; isernia_europolis_2014, )). Deliberation can occur in many different forms, in many different ways, and have many different outcome measures. In what follows, we clarify the aspects of deliberation literature that inform our system.
@@
The abstractions of Plurals map to the practice of deliberation. Ryfe breaks deliberative practice into three phases (ryfe_does_2005, ): (1) The organization of the encounter, (2) the deliberation within the encounter, and (3) the final product of deliberation. Agents are the building blocks of deliberation. As such, Agent initialization corresponds to Phase 1. The deliberation within the encounter is governed by Structures and combination instructions, corresponding to Phase 2. Finally, Moderators can amend the final product of deliberation, corresponding to Phase 3. Separate from Ryfe, Morrell (morrell_listening_2018, ) considers three factors of deliberation that affect outcomes: individual dispositions, institutional structures, and facilitators/moderators. Again, these correspond almost directly to our abstractions of Agents, Structures, and Moderators.
More generally, formal deliberation nowadays often occurs in “mini-publics” (smith_mini-publics_2018, ). These are groups of citizens who come together to deliberate, often in an advisory role. Plurals is an end-to-end generator of simulated social ensembles. This is analogous to reproducing the process of mini-public deliberation.
@@
While the ideals of deliberation are not universally agreed upon, we adopt the dichotomy of “first-generation” and “second-generation” ideals articulated by Bächtiger et al. (bachtiger_deliberative_2018, ). According to Bächtiger et al., the first generation of deliberative theorists (e.g., Habermas (habermas_structural_1991, )) emphasized rationality, achieving a universal consensus, and reason-giving. The second generation of deliberative theorists took a more expansive view of deliberation, beyond rationality and universalism (bachtiger_deliberative_2018, ). For example, second-wave deliberation also valued more emotional forms of communication (neblo_impassioned_2020, ), lived experience, testimony, and storytelling (bachtiger_deliberative_2018, ). Furthermore, to second-wave theorists, the goal was not necessarily a universal consensus (since legitimate disagreement may still exist after perfect deliberation (marti_pluralism_2017, )), but rather a clarifying of understanding (bachtiger_deliberative_2018, ; fraser_rethinking_1990, ).
We incorporate these ideals into our system as both persona templates (how LLMs should enact personas) and combination instructions (how LLMs should combine information with others). To do this, we started with the taxonomy of first-generation and second-generation principles from (bachtiger_deliberative_2018, ). Two authors then engaged in an iterative, two-step process where we first decided whether each dimension was relevant to AI agents, and then how to operationalize this dimension for both generations of deliberation thought. Appendix LABEL:tab:deliberation_ideals lists how we operationalized each ideal.
Some, but not all, ideals or benefits of human deliberation may apply to AI deliberation. Deliberative mini-publics can be useful for the outcomes that they produce (warren_deliberative_1996, ; bachtiger_deliberative_2018, ; smith_mini-publics_2018, ) or the process that produces these outcomes. Regarding the latter, deliberation proponents argue deliberation has certain epistemic (outcome-independent) benefits—such as increased perceived legitimacy for decisions when the sequence of thought leading to them is made public (estlund_epistemic_2018, ). It is the former—outcome-oriented benefits—that is relevant to AI deliberation.
To be clear, our system is inspired by human deliberation; it is not meant to substitute for it. An analogy we offer is how architects and engineers often draw on the natural world to create artifacts. For example, Velcro was inspired by burrs sticking to the inventor’s dog (hook_and_loop_invention_nodate, ). Likewise, multi-agent LLM systems may draw inspiration from the social world—but we do not view LLM deliberation as a substitute for human deliberation. The limits of this metaphor are discussed in section 7.
@@ 2.2. Pluralistic Sociotechnical Systems
Other projects have sought to broaden the representation of technology, what we term “pluralistic sociotechnical systems” for shorthand. These approaches usually focus exclusively on individuals (argyle_out_2023, ), groups (gordon_jury_2022, ; lee_webuildai_2019, ), or governance structures (zhang_policykit_2020, ). As an end-to-end generator of simulated social ensembles, Plurals does all three.
@@
Our system extends prior work aimed at broadening the representation of technological systems through simulation. These approaches address the inherent problems of collapsing diverse viewpoints into a single perspective, a phenomenon we term “output collapse”. In data labeling, annotators often disagree (sandri_why_2023, ; popovic_agree_2021, ; miceli_between_2020, ), yet traditional supervised learning typically resolves these disparities by selecting the majority label. This majority-driven approach can silence minority viewpoints or result in a system that behaves like a “pseudo-human” (gordon_jury_2022, ), presenting a blurred representation that diverges from individual perspectives.
@@
To address output collapse, researchers developed systems that simulate specific perspectives (davani_dealing_2022, ; lee_webuildai_2019, ; he_cura_2023, ; gordon_jury_2022, ). We follow this tradition with Plurals. The most similar system is Juries (gordon_jury_2022, ), which is an architecture and interface for person-specific models. Juries allows end-users to create panels of simulated annotators who make classifications, with the option to upweight dissenting voices.
@@
While the above work primarily addresses individuals or small groups, some systems enhance technology’s representativeness by customizing governance structures. For example, PolicyKit (zhang_policykit_2020, ) allows online communities to create arbitrary governance structures easily, essentially letting communities embed their own values. Similarly, Schneider et al. created the idea of “modular politics” (schneider_modular_2021, ), where communities construct bottom-up governance structures from distinct components.
@@
Large language models (LLMs) have intensified both the problem of output collapse and the potential solutions to combat it. While a single “ground truth” was often contested in conventional classification (sandri_why_2023, ; popovic_agree_2021, ), open-ended text generation further complicates the notion of a single, “correct” answer. Simultaneously, LLMs can potentially be steered to adopt viewpoints through “personas” (jiang_personallm_2024, ; salminen_deus_2024, ; ha_clochat_2024, ). We adopt Argyle et al’s (argyle_out_2023, ) method of generating personas from government datasets to use as LLM prompts. Both Argyle et al.’s method (argyle_out_2023, ) and Gordon et al.’s Juries (gordon_jury_2022, ) employ multiple individual characteristics to construct personas. By using nationally representative datasets, we create personas reflecting general population attributes. These intersectional personas should theoretically enhance diversity beyond single-attribute personas, avoiding ecological fallacies and reducing homogenization (Case Study 1).
@@
Plurals is an evolution and extension of the above ideas. Plurals is an evolution of prior work: What Juries is to aggregative democracy, Plurals is to deliberative democracy (tsai_generative_2024, ). Unlike Juries’ focus on classification labels, Plurals can generate open-ended text. As Gordon et al. (gordon_jury_2022, ) write, “jury learning does not draw on the deliberative nature of juries, which has been the subject of decades of study in legal literature.” This deliberation is our contribution. Plurals also expands the core idea of Juries. In Plurals syntax (Figure 2), a Gordon et al. jury would be an ensemble where Agents complete tasks in parallel without information sharing. This is just one way that multi-agent systems can be designed. By allowing users to create diverse structures and customize Agent deliberation within these structures, we offer a more comprehensive approach to studying and implementing pluralistic AI. Finally, unlike Juries (gordon_jury_2022, ), Plurals does not require a representative dataset with annotator demographics (which can be prohibitive to obtain). By allowing users to change governance structures, Plurals is conceptually similar to PolicyKit. However, Plurals differs from PolicyKit in that Plurals supports the construction of Agents and Moderators (the “before” and “after” of Structures using Ryfe’s three-part terminology of deliberation (ryfe_does_2005, )). In brief, Plurals allows end-to-end generation of simulated social ensembles.

### v6

@@ Abstract.
journalyear: 2025copyright: ccconference: CHI Conference on Human Factors in Computing Systems; April 26-May 1, 2025; Yokohama, Japanbooktitle: CHI Conference on Human Factors in Computing Systems (CHI ’25), April 26-May 1, 2025, Yokohama, Japandoi: 10.1145/3706598.3713675isbn: 979-8-4007-1394-1/25/04ccs: Computing methodologies Artificial intelligenceccs: Computing methodologies Multi-agent systemsccs: Computing methodologies Intelligent agentsccs: Human-centered computing Interaction paradigmsccs: Human-centered computing Interactive systems and toolsccs: Human-centered computing Open source softwareccs: Human-centered computing Interaction design theory, concepts and paradigms
@@ 1. Introduction
Figure 1. System diagram of Plurals—an end-to-end generator of simulated social ensembles. (1) Agents complete tasks within (2) Structures, with communication optionally summarized by (3) Moderators. Plurals integrates with government datasets (1a) and templates inspired by deliberative democracy theory (1b). The building block is Agents, which are large language models (LLMs) that have system instructions and tasks. System instructions can be generated from user input, government datasets (American National Election Studies; ANES), and templates from deliberative democracy literature (Bächtiger et al., 2018). Agents exist within Structures, which define what information is shared. Combination instructions tell Agents how to combine the responses of other Agents when deliberating in the Structure. Users can customize an Agent’s combination instructions or use existing templates drawn from deliberation literature and beyond. Moderators aggregate responses from multi-agent deliberation.
@@
System diagram of Plurals. The system consists of three components: Agents are language models that perform tasks; Structures define information sharing between agents; Moderators summarize agent communication. The visual diagram depicts each of these components and also contrasts the multi-perspective approach of Plurals with the conventional approach of a “view from nowhere”.
There is a fundamental tension between how generative AI models are built and how they are used. Companies typically build a small number of foundation or “generalist” models that dominate the market (Vipra and Korinek, 2023). However, these generalist models are used by a diverse base of users—with varying preferences and values. Invariably, this tension sparked allegations of bias, with supposedly neutral models accused of favoring certain viewpoints (Feng et al., 2023b; Braun and Villasenor, 2023; Fujimoto and Takemoto, 2023).
While a tempting solution is to aim for models that have “no bias” and hold a “view from nowhere” (Haraway, 1997), truly neutral models are likely infeasible. Some scholars argue that all knowledge is situated (Haraway, 1997). But with open-ended text generation, defining some unbiased ground truth is especially difficult. For many use cases, there is no unbiased ground truth. This difficulty is compounded by the fact that users can ask models a large variety of questions. Any bias benchmark can only capture an infinitesimal slice of the query space (Raji et al., 2021).
@@
As an alternative to “bias-free” models, we introduce a new pluralistic AI system (Sorensen et al., 2024), Plurals, that can accomplish these tasks. It is a public-facing Python library (Figure 1 for system overview, Figure 2 for code snippets, see here [^1] https://github.com/josh-ashkinaze/plurals for library). Plurals consists of Agents (optionally integrated with government datasets for nationally representative personas) which deliberate within customizable Structures, with Moderators overseeing deliberation. Plurals is an end-to-end generator of customizable “simulated social ensembles”. We incorporate interaction templates inspired by deliberative democracy theory and integration with government datasets for nationally representative personas. For example, to create an Agent representing a male California resident, our system samples a statistically representative citizen from American National Election Studies, and then uses the citizen’s demographics and political stances as an LLM prompt. We draw on deliberative democracy theory, which emphasizes dialogue between different views (Bächtiger et al., 2018; Martí, 2017), as a blueprint. Our work builds on research in deliberation (Bächtiger et al., 2018; Morrell, 2018; Brown, 2018; Smith and Setälä, 2018; Fraser, 1990; Habermas, 1991; Martí, 2017), pluralistic sociotechnical systems (Argyle et al., 2023b; Lee et al., 2019; Gordon et al., 2022; Zhang et al., 2020), and multi-agent AI alignment approaches (Vahidov and Fazlollahi, 2004; Irving et al., 2018; Talebirad and Nadiri, 2023; Kim et al., 2024; Chan et al., 2023). To our knowledge, Plurals is the first general-purpose toolkit for pluralistic, multi-agent interactions modeled after deliberative democracy.
@@
  • Theoretical: We created a multi-agent system incorporating ideals of deliberative democracy theory. Our system also introduces “interactional pluralism”, a pluralism that exists not only in the distribution of agent properties but also in the protocols governing their interactions. Users can customize how Agents should combine information with each other and the information-sharing structures in which Agents exist.
@@ 1.1. Brief System Overview
Plurals allows users to create simulated social ensembles with Agents, Structures, and Moderators: Agents complete tasks within Structures, which define how information is shared between Agents. Moderators can summarize multi-agent communication. Each abstraction is highly customizable. Agents can use various LLMs and have system instructions set manually, through persona generation, or via American National Election Studies (ANES) integration. Structures vary in information-sharing, complexity, and randomness. For example, users can define custom networks of Agents in a few lines of code (Figure 2). The behavior of Agents within Structures (how they should combine information from other agents) can be tuned via combination instructions. Our package comes pre-populated with templates for personas, combination instructions, and moderators—drawing on deliberative democracy theory and prior work.
@@ 2. System Grounding
Plurals is grounded in deliberation literature, sociotechnical systems that broaden technological perspectives, and multi-agent systems for AI alignment. The result is an end-to-end generator of simulated social ensembles—groups that engage in deliberation. We integrate deliberative theory into our system by incorporating templates of first- and second-generation deliberative ideals and using deliberative theory to inform the structure of AI deliberation. We extend previous work on broadening technological perspectives, such as Argyle et al.’s dataset-based personas (Argyle et al., 2023b), Gordon et al.’s “juries” (Gordon et al., 2022), and Zhang et al.’s PolicyKit (Zhang et al., 2020). Our system encompasses individual, group, and governance-level simulations, unlike previous approaches that focused on flexibility at only one of these three levels. By drawing on the concept of deliberative “mini-publics” (groups who engage in deliberation (Smith and Setälä, 2018)), we evolve from aggregative methods (like juries) to a more deliberative approach. Additionally, we contribute to multi-agent AI research by offering a flexible system for creating diverse interaction structures and providing a reusable infrastructure for experiments.
@@ 2.1. Deliberation
Deliberation is defined as “mutual communication that involves weighing and reflecting on preferences, values, and interests regarding matters of common concern” (Bächtiger et al., 2018). As Bächtiger et al. distinguish (Bächtiger et al., 2018), deliberative democracy differs from aggregative democracy. The former centers talking and the latter centers voting—though they can co-occur (e.g., talking before voting (Fishkin, 2003; Isernia and Fishkin, 2014)). Deliberation occurs in many different forms, in many different ways, and has many different outcome measures. In what follows, we clarify the aspects of deliberation literature that inform our system.
@@
The abstractions of Plurals map to the practice of deliberation. Ryfe breaks deliberative practice into three phases (Ryfe, 2005): (1) The organization of the encounter, (2) the deliberation within the encounter, and (3) the final product of deliberation. Agents are the building blocks of deliberation. As such, Agent initialization corresponds to Phase 1. The deliberation within the encounter is governed by Structures and combination instructions, corresponding to Phase 2. Finally, Moderators can amend the final product of deliberation, corresponding to Phase 3. Separate from Ryfe, Morrell (Morrell, 2018) considers three factors of deliberation that affect outcomes: individual dispositions, institutional structures, and facilitators/moderators. Again, these correspond almost directly to our abstractions of Agents, Structures, and Moderators. More generally, formal deliberation nowadays often occurs in “mini-publics” (Smith and Setälä, 2018). These are groups of citizens who come together to deliberate, often in an advisory role. Plurals is an end-to-end generator of simulated social ensembles. This is analogous to reproducing the process of mini-public deliberation.
@@
While the ideals of deliberation are not universally agreed upon, we adopt the dichotomy of “first-generation” and “second-generation” ideals articulated by Bächtiger et al. (Bächtiger et al., 2018). According to Bächtiger et al., the first generation of deliberative theorists (e.g., Habermas (Habermas, 1991)) emphasized rationality, achieving a universal consensus, and reason-giving. The second generation of deliberative theorists took a more expansive view of deliberation, beyond rationality and universalism (Bächtiger et al., 2018). For example, second-wave deliberation also valued more emotional forms of communication (Neblo, 2020), lived experience, testimony, and storytelling (Bächtiger et al., 2018). Furthermore, to second-wave theorists, the goal was not necessarily a universal consensus (since legitimate disagreement may still exist after perfect deliberation (Martí, 2017)), but rather a clarifying of understanding (Bächtiger et al., 2018; Fraser, 1990).
We incorporate these ideals into our system as both persona templates (how LLMs should enact personas) and combination instructions (how LLMs should combine information with others). To do this, we started with the taxonomy of first-generation and second-generation principles from (Bächtiger et al., 2018). Two authors then engaged in an iterative, two-step process where we first decided whether each dimension was relevant to AI agents, and then how to operationalize this dimension for both generations of deliberation thought. Appendix LABEL:tab:deliberation_ideals lists how we operationalized each ideal.
Some, but not all, ideals or benefits of human deliberation may apply to AI deliberation. Deliberative mini-publics can be useful for the outcomes that they produce (Warren, 1996; Bächtiger et al., 2018; Smith and Setälä, 2018) or the process that produces these outcomes. Regarding the latter, deliberation proponents argue deliberation has certain epistemic (outcome-independent) benefits—such as increased perceived legitimacy for decisions when the sequence of thought leading to them is made public (Estlund and Landemore, 2018). It is the former—outcome-oriented benefits—that is relevant to AI deliberation.
To be clear, our system is inspired by human deliberation; it is not meant to substitute for it. By analogy, engineers often draw on the natural world to create artifacts. For example, Velcro was inspired by burrs sticking to the inventor’s dog (Loop, [n. d.]). The limits of human deliberation as a metaphor are discussed in section 7.
@@ 2.2. Pluralistic Sociotechnical Systems
Other projects have sought to broaden the representation of technology, what we term “pluralistic sociotechnical systems” for shorthand. These approaches usually focus exclusively on individuals (Argyle et al., 2023b), groups (Gordon et al., 2022; Lee et al., 2019), or governance structures (Zhang et al., 2020). As an end-to-end generator of simulated social ensembles, Plurals does all three.
@@
Our system extends prior work aimed at broadening the representation of technological systems through simulation. These approaches address the inherent problems of collapsing diverse viewpoints into a single perspective, a phenomenon we term “output collapse”. In data labeling, annotators often disagree (Sandri et al., 2023; Popović, 2021; Miceli et al., 2020), yet traditional supervised learning typically resolves these disparities by selecting the majority label. This majority-driven approach can silence minority viewpoints or result in a system that behaves like a “pseudo-human” (Gordon et al., 2022), presenting a blurred representation that diverges from individual perspectives.
@@
To address output collapse, researchers developed systems that simulate specific perspectives (Davani et al., 2022; Lee et al., 2019; He et al., 2023; Gordon et al., 2022). Plurals follows this tradition. The most similar system is Juries (Gordon et al., 2022), which is an architecture and interface for person-specific models. Juries allows end-users to create panels of simulated annotators who make classifications, with the option to upweight dissenting voices.
@@
While the above work primarily addresses individuals or small groups, some systems enhance technology’s representativeness by customizing governance structures. For example, PolicyKit (Zhang et al., 2020) allows online communities to create arbitrary governance structures easily, essentially letting communities embed their own values. Similarly, Schneider et al. created “modular politics” (Schneider et al., 2021), where communities construct governance structures from distinct components.
@@
Large language models (LLMs) have intensified both the problem of output collapse and the potential solutions to combat it. While a single “ground truth” was often contested in conventional classification (Sandri et al., 2023; Popović, 2021), open-ended text generation further complicates the notion of a single, “correct” answer. Simultaneously, LLMs can potentially be steered to adopt viewpoints through “personas” (Jiang et al., 2024; Salminen et al., 2024; Ha et al., 2024). We adopt Argyle et al.’s (Argyle et al., 2023b) method of generating personas from government datasets to use as LLM prompts. Both Argyle et al.’s method (Argyle et al., 2023b) and Gordon et al.’s Juries (Gordon et al., 2022) employ multiple individual characteristics to construct personas. By using nationally representative datasets, we create personas reflecting general population attributes. These intersectional personas should theoretically enhance diversity beyond single-attribute personas, reducing homogenization (Case Study 1).
@@
Plurals is an evolution and extension of the above ideas. Plurals is an evolution of prior work: What Juries is to aggregative democracy, Plurals is to deliberative democracy (Tsai et al., 2024). Unlike Juries’ focus on classification labels, Plurals can generate open-ended text. As Gordon et al. (Gordon et al., 2022) write, “jury learning does not draw on the deliberative nature of juries, which has been the subject of decades of study in legal literature.” This deliberation is our contribution. Plurals also expands the core idea of Juries. In Plurals syntax (Figure 2), a Gordon et al. jury is an ensemble where Agents complete tasks in parallel without information sharing. This is just one communication structure. By allowing users to create diverse structures and customize Agent deliberation within these structures, we offer a more comprehensive approach to studying and implementing pluralistic AI. Finally, unlike Juries (Gordon et al., 2022), Plurals does not require a task-specific representative dataset with annotator demographics (which can be prohibitive to obtain). By allowing users to change governance structures, Plurals is conceptually similar to PolicyKit. However, Plurals differs from PolicyKit in that Plurals supports the construction of Agents and Moderators (the “before” and “after” of Structures using Ryfe’s three-part terminology of deliberation (Ryfe, 2005)). In brief, Plurals allows end-to-end generation of simulated social ensembles.